# Make Me Happier: Evoking Emotions Through Image Diffusion Models

### v3

@@ Abstract
Despite the rapid progress in image generation, emotional image editing remains under-explored. The semantics, context, and structure of an image can evoke emotional responses, making emotional image editing techniques valuable for various real-world applications, including treatment of psychological disorders, commercialization of products, and artistic design. For the first time, we present a novel challenge of emotion-evoked image generation, aiming to synthesize images that evoke target emotions while retaining the semantics and structures of the original scenes. To address this challenge, we propose a diffusion model capable of effectively understanding and editing source images to convey desired emotions and sentiments. Moreover, due to the lack of emotion editing datasets, we provide a unique dataset consisting of 340,000 pairs of images and their emotion annotations. Furthermore, we conduct human psychophysics experiments and introduce four new evaluation metrics to systematically benchmark all the methods. Experimental results demonstrate that our method surpasses all competitive baselines. Our diffusion model is capable of identifying emotional cues from original images, editing images that elicit desired emotions, and meanwhile, preserving the semantic structure of the original images. All code, model, and dataset will be made public.
Figure 1: The generated images evoke a sense of happiness in viewers, contrasting with the negative emotions elicited by the source images. Given a source image that triggers negative emotions (framed in green), our method (Ours) synthesizes a new image that elicits the given positive target emotions (in red), while maintaining the essential elements and structures of the scene. For instance, in the first quadrant where trash on the ground evokes negative emotions, our model performs image editing by eliminating the trash while preserving the outdoor context, including trees and grass, in the same spatial arrangement as the original scene. For comparisons, we include other competitive methods (Sec.4). Zoom in for better views.
@@ 1 Introduction
“I am feeling down. Enhance my room to bring in more excitement."
What we perceive not only entails useful visual information but also evokes profound emotional reactions. Incorporating visual cues that elicit emotions into images holds practical significance in various domains. For instance, in commercials, it amplifies product branding and shapes public opinions, captivating consumer attention. Similarly, integrating emotion-evoking elements into AR/VR experiences aids in addressing psychological disorders such as autism and schizophrenia. Despite significant advancements in image generation techniques Gal et al. (2022); Wang et al. (2023); Yu et al. (2018), there has been a notable absence of focus given to emotion conditioning. Here, we introduce the novel and important problem of emotion-evoked image generation — given an initial image and a target emotion, the task is to create an image that elicits the specified emotion of human viewers while maintaining the original scene’s semantics and structures.
Emotion-evoked image generation poses a challenge as it requires a deep understanding and identification of the subtle context and semantic elements present in the source images that evoke emotional responses. Additionally, this task also requires image editing capabilities to effectively remove, replace, add, or modify elements within the source images to elicit the intended emotional response. For instance, in Figure 1, the presence of a burning lawn significantly contributes to feelings of anger. Consequently, in the generated image, the burning lawn is substituted with a serene scene featuring a cluster of flowers, thereby transforming the overall emotional response to awe.
Studies Yang et al. (2023, 2018) have identified both global and local factors influencing emotional states, such as overall color and tonal shifts in image backgrounds, local facial expressions, and the presence of emotion-associated objects like graveyards correlated with sadness and balloons correlated with amusement. As also seen in Figure 2, introducing fearful elements like fire or a woman in white into forests (Column 2) can readily evoke negative emotions, as these elements prompt instinctual responses to unpleasant or adverse environments. However, not all concepts serve as emotional triggers; some may remain neutral, such as the sky, plants, or windows. In the same examples in Column 2, forests are neutral concepts and do not elicit emotional responses. Given that human emotions tend to exhibit greater sensitivity to negative stimuli over positive stimuli Baumeister et al. (2001); Ito et al. (1998), simply manipulating global intensity by adjusting the brightness of entire images in Column 2 would not necessarily alter the negative emotions.
@@
Drawing inspiration from these emotion-associated factors, we propose our model, EmoEditor, for emotion-evoked image generation. We base EmoEditor on state-of-the-art stable diffusion models and introduce three technical novelties. (1) EmoEditor employs a novel dual-branch architecture in the reverse diffusion process, that integrates emotion-conditioned global context and local emotional cues from the source images, to ensure coherence with the original content while reflecting the desired emotional outcome. (2) During training, the network’s behaviors are guided by aligning model creativity with human expectations in the neuro-symbolic space, implicitly learning the mapping from target emotions to human-annotated text instructions. (3) During inference, iterative emotion discrimination mechanisms are employed to autonomously select emotionally coherent images.
Before the introduction of our EmoEditor, traditional methods, such as color transfer and style transfer Pitié et al. (2007); Gatys et al. (2015); Weng et al. (2023), were commonly employed to modify the emotional tone of images. These methods typically require reference emotional images or textual cues as additional inputs. However, they face challenges in achieving satisfactory performance in eliciting desired emotions because they struggle to understand, identify, and manipulate the local regions of source images that evoke emotional responses. In Figure 2, while darkening brightness or deepening hues may easily evoke fear (Example 1), simply brightening colors in an image of a forest fire is inadequate for eliciting amusement (Example 4).
@@
Figure 2: Emotions are influenced by both global and local cues from the source images. We present two image generation examples showcasing the transition from positive to negative emotions in Column 1 (shaded in blue) and two examples from negative to positive emotions in Column 2 (shaded in orange). In each example, the target emotions are in red, and source images are framed in green. Generated results from three methods are presented. Zoom in for better views.
@@
With the rapid advancements in diffusion models, various effective text-to-image editing techniques have emerged Kawar et al. (2023); Brooks et al. (2023); Ruiz et al. (2023). However, these methods are not specifically tailored for understanding and identifying emotion-associated elements within images. To address this gap, we propose a competitive baseline (Large Model Series), concatenating multiple large language and vision models. This baseline first understands source images through image captioning, then generates emotion-prompted instructions using language models, and finally edits images prompted by these instructions with diffusion models. Upon comparison with existing methods through a series of human psychophysics experiments and four new quantitative evaluations, our end-to-end trained EmoEditor achieves outstanding image generation results. It effectively preserves the structural coherence and semantic consistency wtih the source images while eliciting the target emotion of human viewers.
To obtain training data for emotion-evoked image generation, we establish a dataset curation pipeline and introduce the first large-scale EmoPair dataset, comprising approximately 340,000 pairs of images annotated with emotion-conditioned editing instructions. Unlike existing datasets on visual emotion analysis Mikels et al. (2005); Machajdik, Hanbury (2010); You et al. (2016); Peng et al. (2015); Yang et al. (2023), which assign emotion labels to individual images for classification, our dataset provides paired images annotated with source and target emotion labels.
@@
The main contributions of this paper are highlighted below: (1) We introduce a new and important problem of emotion-evoked image generation. To benchmark all methods, we establish a comprehensive framework for evaluating model performances, incorporating human psychophysics experiments and four newly introduced metrics. (2) We introduce EmoEditor, an emotion-evoked diffusion model, which employs a novel two-branch architecture to integrate global context with local semantic regions of source images that evoke emotional responses. (3) EmoEditor learns emotional cues during end-to-end training, eliminating the need for additional emotion reference images. During inference, EmoEditor generates emotion-evoked images without hand-crafted emotion-conditioned text instructions, while preserving context and structural coherence with the source image. Its performance is superior among all the competitive methods. (4) We curate the first large-scale EmoPair dataset, containing 340,000 image pairs with emotion annotations.
@@ 2.1 Visual Emotion Analysis
Visual Emotion Analysis (VEA) aims to understand and predict human emotional responses from visual data. In psychology, Categorical Emotion States (CES) form the basis for emotion classification models like the 8-category Mikels model Mikels et al. (2005). Early research used manually crafted features such as color and content Machajdik, Hanbury (2010), while recent advancements employ deep learning to capture emotional cues more accurately You et al. (2015); Yang et al. (2018). Despite progress, generating images to evoke specific emotions remains underexplored. We propose EmoEditor, a two-branch diffusion model that integrates emotion recognition and image generation to address this. Unlike the Affective Image Filter Weng et al. (2023), which requires explicit textual instructions, EmoEditor does not need text guidance to achieve desired emotions.
Existing VEA datasets Peng et al. (2015); You et al. (2016) focus on emotion classification, including EmoSet Yang et al. (2023), which has 120k images with emotion labels and attributes. However, they lack image pairs with annotated emotions and text editing instructions essential for emotion-evoked image generation. Thus, we introduce the EmoPair dataset, comprising 340k image pairs with differing source and target emotions while ensuring consistent scene semantics and structure.
@@ 2.2 Image Editing
Color and Style Transfer. Methods like PDF-TransferPitié et al. (2007), Neural Style TransferGatys et al. (2015), and CLIP-StylerKwon, Ye (2022) manipulate images using reference images, but they struggle with local region adjustments, limiting their versatility in generating emotion-evoking images. Our EmoEditor balances local and global manipulations without needing additional references.
Text-to-Image Editing. Text-to-image editing is increasingly popular. Earlier studies used GANs Goodfellow et al. (2014) and CLIP models Radford et al. (2021) for image manipulation based on detailed text descriptions Patashnik et al. (2021); Abdal et al. (2022). Recent diffusion models like Imagic Kawar et al. (2023), Prompt-to-prompt Hertz et al. (2022), InstructPix2Pix Brooks et al. (2023), and DreamBooth Ruiz et al. (2023) generate high-quality images from text prompts but often ignore emotional aspects and rely heavily on detailed instructions. In contrast, our EmoEditor needs only target emotion and source image inputs, using end-to-end training to understand and edit emotional cues while preserving image structure and coherence.
@@ 3 Method
We formulate emotion-evoked image generation as a supervised learning task. First, we curate the EmoPair dataset containing image pairs with emotion labels and corresponding text instructions for editing (Sec.3.1). Next, EmoEditor is trained on this dataset (Sec.3.2). During inference, it takes source images and target emotions, producing generated images without the need for human-annotated text instructions or emotion reference images.
Figure 3: Pipeline for Curating Our EmoPair Dataset. The dataset comprises two subsets: EmoPair-Annotated Subset (EPAS, left blue box) and EmoPair-Generated Subset (EPGS, right orange box). Each subset includes schematics depicting the creation, selection, and labeling of image pairs in the upper quadrants, with two example pairs in the lower quadrants. Each example pair comprises a source image (framed in green) and a target image. The classified source and target emotion labels (highlighted in red) and target-emotion-driven text instructions for image editing are provided. See more examples in Supp.Sec.S1. See Sec.3.1 for EmoPair dataset.
@@ 3.1 Curating Our EmoPair Dataset
Given the unavailability of a paired image dataset for emotion editing, we curate EmoPair. It consists of two subsets: EmoPair-Annotated Subset (EPAS), containing 331,595 image pairs from Ip2pBrooks et al. (2023) annotated with emotion labels; and EmoPair-Generated Subset (EPGS), consisting of 6,949 pairs generated based on text instructions given target emotions. All images from EmoPair are of size 224$\times$224. Figure 3 shows the pipeline for curating our EmoPair dataset.
Emotion Predictor $\mathcal{P}$. To better construct our dataset, we first train an emotion predictor $\mathcal{P}$ on the existing VEA dataset EmoSet Yang et al. (2023). All images are classified into 8 emotion categories, encompassing four positive (amusement, awe, contentment, excitement) and four negative (anger, disgust, fear, sadness) emotions. $\mathcal{P}$ employs ResNet18 He et al. (2016) as a backbone and achieves a top-1 accuracy of 73% on EmoSet’s test set.
@@
EmoPair-Generated Subset (EPGS). To address the lack of emotional cues in the Ip2p dataset, we create the EPGS subset using emotionally rich images from EmoSet Yang et al. (2023). However, EmoSet lacks paired images with consistent scenes and target emotions, and it doesn’t include text editing instructions. Single-word prompts like “[Target Emotion]" prove ineffective in existing text-to-image models due to their limited emotional reasoning. Single-emotion terms lack specificity, resulting in varied interpretations depending on the context of source images. For instance, enhancing an image for excitement can vary significantly based on context, such as prompting a happy smile for a person or adding fireworks to the sky for a landscape. Thus, to generate accurate image pairs, we develop 50 general instructions for transitioning to desired emotions of 8 categories using GPT-3 Brown et al. (2020). Human annotators rank these instructions, keeping the top ten per emotion. The pre-trained Ip2p model then uses these instructions to manipulate EmoSet source images, resulting in 6,949 high-quality image pairs after quality control measures (see Supp.Sec.S1).
@@ 3.2 Our Diffusion Model - EmoEditor
Figure 4: Architecture of Our Proposed EmoEditor. EmoEditor is an image diffusion model, consisting of local (shaded in green) and global (shaded in blue) branches. The pre-trained VAE’s encoder $\mathcal{E}$ and decoder $\mathcal{D}$ remain fixed throughout training and inference. Exclusively employed during inference, the fixed emotion predictor $\mathcal{P}$ predicts the emotions on generated images for the iterative emotion inference. See Sec.3 for details.
@@
Our EmoEditor, in Figure 4, takes a source image $I_{s}$ and a target emotion $y$ to generate an image evoking the desired emotion. It employs the latent diffusion model (LDM) Rombach et al. (2022) with three key novelties. EmoEditor avoids using emotion labels for $I_{s}$ for two reasons: (1) Obtaining emotion labels for source images is often impractical in real-world scenarios. (2) Source images can be complex, conveying multiple emotions simultaneously. Utilizing a stereotypical source emotion as input potentially introduces biases and limits the model’s ability to generate diverse outputs.
@@
Integration of Global and Local Emotion Cues. In the global branch of our EmoEditor, we convert the target emotion $y$ into a binary one-hot vector $e_{oh}$, where the corresponding entry to $y$ in $e_{oh}$ is set to 1 and the rest are set to 0. This simplifies model design, enhancing comprehension and generation of complex emotional images, as one-hot encoding can capture intricate emotion combinations through a multi-label representation. Conversely, text inputs demand extra transformation steps, potentially introducing noise or errors. We introduce an emotion encoder $\tau_{\theta}$ to encode $e_{oh}$ into $e$ via fully connected layers: $e=\tau_{\theta}(e_{oh})$ (see Supp.Sec.S2.1 for details), trained from scratch to learn neuro-symbolic embeddings. Later on, we introduce novel losses to regularize the behaviors of $\tau_{\theta}$. Notably, we intentionally avoid using explicit text editing instructions as conditional inputs to the model due to the varied interpretations of target emotions, where multiple solutions can elicit the same emotions. Including explicit text instructions may hinder the model’s creativity, restricting its capacity to produce diverse sets of images conveying the same target emotions.
@@
In the local branch of our EmoEditor, we introduce an image encoder $\mathcal{E}$ to extract visual features of the source image, acquiring $z$ for $I_{s}$: $z=\mathcal{E}(I_{s})$. Similar to the LDM Rombach et al. (2022), $\mathcal{E}$ is based on a pre-trained VAEKingma, Welling (2022). The forward diffusion process gradually adds noise to $z$, generating a noisy latent $z_{t}$, where the noise level increases over $t$ time steps ($t\in T$). Specifically, we expressed $z_{t}=\sqrt{\alpha_{t}}\cdot z_{t-1}+\sqrt{1-\alpha_{t}}\cdot\epsilon$, where $\alpha_{t}$ is the diffusion coefficient controlling the rate of noise increase, and $\epsilon$ is a random noise sampled from a normal distribution.
@@
Importantly, we differentiate our alignment loss from the losses typically employed in classical language models, which focus on predicting specific word tokens. Our approach loosens the constraint of precisely predicting the exact text instructions. This flexibility empowers our model to explore a wider array of image editing solutions, that can augment the source emotions of human viewers to target emotions. Consequently, our model learns to reason in a neuro-symbolic manner akin to humans, and meanwhile, preserves its creativity and capacity for exploring novel editing solutions.
Iterative Emotion Inference. Editing emotions is a complex process often requiring iterative image edits. To address this, we introduce a recurrent emotion critic process during inference. This critic iteratively infers the evoked emotions from generated images while assessing their structural coherence and semantic consistency. Specifically, we employ the emotion predictor $\mathcal{P}$ as the critic.
In the first iteration of image generation, the source image $I_{s}$ serves as both the input and condition image, producing latent variables $z$ and $z_{t}$. These, along with the emotion vector $e$, are fed to the EmoEditor to obtain the generated image $\hat{I}$. The critic evaluates the generated image quality based on two criteria: (1) the structural similarity (SSIM) between $\hat{I}$ and $I_{s}$ falls within 0.3-0.8, and (2) the predicted emotion by $\mathcal{P}$ on $\hat{I}$ matches the target emotion, with a confidence level exceeding 0.8. If both criteria are met, EmoEditor ceases image generation; otherwise, it continues to generate new images. In subsequent iterations, we use $\hat{I}$ as the input image, generating its noisy latent variable $\hat{z}_{t}$, with $I_{s}$ remaining as the condition. Along with the emotion vector $e$, the EmoEditor generates new edited images for the critic. This process iterates until both criteria are satisfied, or the number of critic iterations exceeds a predefined limit of 30. In the latter case, we select the generated result with the highest confidence level predicted by $\mathcal{P}$ given the target emotion among all past iterations.
See Supp.Sec.S2.2 for Implementation Details.

### v4

@@ Abstract
Despite the rapid progress in image generation, emotional image editing remains under-explored. The semantics, context, and structure of an image can evoke emotional responses, making emotional image editing techniques valuable for various real-world applications, including treatment of psychological disorders, commercialization of products, and artistic design. First, we present a novel challenge of emotion-evoked image generation, aiming to synthesize images that evoke target emotions while retaining the semantics and structures of the original scenes. To address this challenge, we propose a diffusion model capable of effectively understanding and editing source images to convey desired emotions and sentiments. Moreover, due to the lack of emotion editing datasets, we provide a unique dataset consisting of 340,000 pairs of images and their emotion annotations. Furthermore, we conduct human psychophysics experiments and introduce a new evaluation metric to systematically benchmark all the methods. Experimental results demonstrate that our method surpasses all competitive baselines. Our diffusion model is capable of identifying emotional cues from original images, editing images that elicit desired emotions, and meanwhile, preserving the semantic structure of the original images. All code, model, and dataset are available at GitHub.
Figure 1: The generated images evoke a sense of happiness in viewers, contrasting with the negative emotions elicited by the source images. Given a source image that triggers negative emotions (framed in green), our method (Ours) synthesizes a new image that elicits the given positive target emotions (in red), while maintaining the essential elements and structures of the scene. For instance, our method replaces the anger-inducing flames in the source image with cute-shaped lamps to evoke the target emotion of amusement. While in an outdoor setting, the raging fire is substituted with a tranquil, lush meadow to inspire a sense of awe. For comparisons, we include other competitive methods. The blue number below each image represents its CAM-based ESMI score, with higher values being better.
@@ 1 Introduction
“I am feeling down. Bring some excitement to my room."
What we perceive not only entails useful visual information but also evokes profound emotional reactions. Incorporating visual cues that elicit emotions into images holds practical significance in various domains. In commercialization, it enhances brand impact and captures consumer attention. In augmented or virtual reality, it supports therapeutic interventions for conditions such as autism and schizophrenia. Despite recent advances in image generation [5, 27, 37], emotion conditioning remains largely underexplored. We introduce the novel task of emotion-evoked image generation: given a source image and a target emotion, the goal is to generate a new image that evokes the specified emotion while preserving the semantics and structure of the original scene. This task is challenging, requiring both recognition of subtle emotional cues in the source image and the ability to modify them meaningfully. For example, as shown in Figure 1, clusters of fire flames on the water intensify the emotion of anger. In the generated image, they are replaced with lamps designed in cute shapes, transforming the overall emotional response to amusement.
Studies have shown that both global and local factors influence emotional responses in images, such as overall color tones, local facial expressions, and emotion-related objects like graveyards (sadness) or balloons (amusement) [33, 32]. As shown in Figure 2, inserting fearful elements, such as fire, into a peaceful scene can evoke negative emotions, as they trigger instinctive responses to threats. However, some elements, such as the sky and plants, remain emotionally neutral. Given that humans are generally more sensitive to negative than positive stimuli [2, 10], simply adjusting global attributes like brightness may not be enough to alter negative emotions. Traditional methods such as color transfer and style transfer [20, 6, 30] attempt to modify emotional tone through global appearance changes. However, they often fail to elicit desired emotions, as they cannot identify and edit the specific local regions responsible for emotional impact. In Figure 2, while darkening the scene may enhance fear (left panel), brightening a scene of a burning hillside does not effectively convey awe (right panel). In contrast, transforming anger into awe requires editing key emotional cues, for example, removing the fire and adding vibrant clouds to create a tranquil atmosphere. Emotion-evoked image generation thus requires not just global color or style adjustments, but also semantic understanding and purposeful local modification.
@@
Drawing inspiration from these emotion-associated factors, we propose our model, EmoEditor, for emotion-evoked image generation. We develop EmoEditor based on state-of-the-art stable diffusion models and introduce three technical novelties. (1) EmoEditor employs a novel dual-branch architecture in the reverse diffusion process that integrates emotion-conditioned global context and local emotional cues from the source images to ensure coherence with the original content while reflecting the desired emotional outcome. (2) During training, the network’s behaviors are guided by aligning model creativity with human expectations in the neuro-symbolic space, implicitly learning the mapping from target emotions to human-annotated text instructions. (3) During inference, iterative emotion discrimination mechanisms are employed to select emotionally coherent images autonomously.
@@
Figure 2: Emotions are influenced by both global and local cues from the source images. We present examples of transitions from positive to negative emotions (blue shading) and from negative to positive emotions (orange shading). In each example, the target emotions are in red, and source images are framed in green.
@@
With the rapid advancements in diffusion models, various effective image editing techniques have emerged [11, 3, 23]. However, these methods are not specifically tailored for understanding and identifying emotion-associated elements within images. To address this gap, we propose a competitive baseline called Large Model Series (LMS), concatenating multiple large language and vision models. This baseline first understands source images through image captioning, then generates emotion-prompted instructions using language models, and finally edits images prompted by these instructions with diffusion models. Upon comparison with existing methods through a series of human psychophysics experiments and a new quantitative evaluation, our end-to-end trained EmoEditor achieves outstanding image generation results. It effectively preserves the structural coherence and semantic consistency with the source images while eliciting the target emotion of human viewers. To obtain training data for this task, we establish a dataset curation pipeline and introduce the first large-scale EmoPair dataset, comprising approximately 340,000 pairs of images annotated with emotion-conditioned editing instructions. Unlike existing datasets on visual emotion analysis [17, 19, 33], which assign emotion labels to individual images for classification, our dataset provides paired images annotated with source and target emotion labels.
@@
The main contributions of this paper are highlighted below: (1) We study the problem of emotion-evoked image generation, which focuses on modifying visual content to trigger intended emotional responses while preserving scene structures and contextual coherence. To benchmark all methods, we establish a comprehensive framework for evaluating model performance, incorporating human psychophysics experiments and introducing a new metric. (2) We propose EmoEditor, an emotion-evoked diffusion model, which employs a novel two-branch architecture to integrate global context with local semantic regions of source images that evoke emotional responses. (3) EmoEditor learns emotional cues during end-to-end training, eliminating the need for additional emotion reference images. During inference, EmoEditor generates emotion-evoked images without hand-crafted emotion-conditioned text instructions, while preserving context and structural coherence with the source image. Its performance is superior among all the competitive methods. (4) We curate the large-scale EmoPair dataset, containing 340,000 image pairs with emotion annotations.
@@ 2.1 Visual Emotion Analysis
Visual Emotion Analysis (VEA) seeks to understand and predict human emotional responses from visual data. In psychology, Categorical Emotion States (CES) underpin classification models such as the 8-category Mikels model [17]. Early work relied on handcrafted features like color [15], while recent methods leverage deep learning for more accurate emotion prediction [35, 32]. However, generating images to evoke specific emotions remains underexplored. To address this, we introduce EmoEditor, a two-branch diffusion model that jointly learns emotion recognition and image generation.
Existing VEA datasets [19, 36] focus on emotion classification, including EmoSet [33], which has 120k images with emotion labels and attributes. However, they lack image pairs with annotated emotions and text editing instructions essential for emotion-evoked image generation. Thus, we introduce the EmoPair dataset, comprising 340k image pairs with differing source and target emotions while ensuring consistent scene semantics and structure.
@@ 2.2 Image Editing
Diffusion-based Image Editing. Earlier studies used GANs [7] and CLIP models [21] for image manipulation based on detailed text descriptions [1]. Diffusion models have advanced, generating high-quality images from various prompts [8, 23]. Methods like SDEdit [16] use stochastic differential equations for editing. Ip2p [3] allows image editing based on human instructions. PnP [25] achieves fine-grained control of generated structures by manipulating spatial features and self-attention. CtrlNet [38] adds conditional control for better precision. Despite these advances, current diffusion models often overlook emotional cues needed for effective emotion editing.
Emotion Editing. Color and style transfer methods like PDF-Transfer [20], Neural Style Transfer (NST) [6], and CLIP-Styler (Csty) [13] manipulate images using reference images but struggle with local adjustments, limiting their versatility in generating emotion-evoking images. AIF [30] reflects emotions from text to images by incorporating affective content into the image generation process. It relies heavily on detailed textual prompts and cannot edit image regions locally. EmoEdit [34] maps input images and target emotions into an embedding guided by a set of fixed learned queries, restricting emotional cues and diversity in generation to the representational capacity of these queries. In contrast, our EmoEditor uses end-to-end training to implicitly understand and edit emotional cues locally and globally without relying on external references.
@@ 3 Method
We formulate emotion-evoked image generation as a supervised learning task. First, we curate the EmoPair dataset with emotion-labeled image pairs and editing instructions. Next, EmoEditor is trained on this dataset. During inference, it takes source images and target emotions, producing generated images without the need for text instructions or emotion reference images.
@@ 3.1 Curating Our EmoPair Dataset
Given the unavailability of a paired image dataset for emotion editing, we curate EmoPair. It consists of two subsets: EmoPair-Annotated Subset (EPAS), containing 331,595 image pairs from Ip2p [3] annotated with emotion labels; and EmoPair-Generated Subset (EPGS), consisting of 6,949 pairs generated based on text instructions given target emotions. All images from EmoPair are of size 224$\times$224. Figure 3 illustrates the pipeline for curating our EmoPair dataset. See Sec S1 for more details.
Emotion Predictor $\mathcal{P}$. To construct our dataset, we use the emotion predictor $\mathcal{P}$ proposed in [33], which was trained on the largest VEA dataset, EmoSet. All images are classified into 8 emotion categories, encompassing four positive (amusement, awe, contentment, excitement) and four negative (anger, disgust, fear, sadness) emotions.
Figure 3: Pipeline for Curating Our EmoPair Dataset. The dataset comprises two subsets: EmoPair-Annotated Subset (EPAS, left blue box) and EmoPair-Generated Subset (EPGS, right orange box). Each subset includes schematics depicting the creation, selection, and labeling of image pairs in the upper quadrants, with two example pairs in the lower quadrants. Each example pair comprises a source image (framed in green) and a target image. The classified source and target emotion labels (highlighted in red) and target-emotion-driven text instructions for image editing are provided. Note that these instructions are used during model training to guide the learning of rich neural-symbolic embeddings, but are not used during model inference.
@@
EmoPair-Generated Subset (EPGS). To address the lack of emotional cues in the Ip2p dataset, we create the EPGS subset using emotionally rich images from EmoSet [33]. However, EmoSet lacks paired images with consistent scenes and target emotions, and it does not include text editing instructions. Single-word prompts like “[Target Emotion]" prove ineffective in existing text-to-image models due to their limited emotional reasoning. Single-emotion terms lack specificity, resulting in varied interpretations depending on the context of source images. For instance, enhancing an image for excitement can vary significantly based on context, such as having a happy smile on a person or adding fireworks to the sky for a landscape. Thus, to generate accurate image pairs, we develop 50 general instructions for transitioning to desired emotions of 8 categories using GPT-3 [4]. Human annotators rank these instructions, keeping the top ten per emotion. The pre-trained Ip2p model then uses these instructions to manipulate EmoSet source images, resulting in 6,949 high-quality image pairs after quality control measures. To ensure the quality of the generated images, we have human annotators providing corrections for some results (see Sec S1).
@@ 3.2 Our Diffusion Model - EmoEditor
Figure 4: Architecture of Our Proposed EmoEditor. EmoEditor is an image diffusion model with local (shaded in green) and global (shaded in blue) branches. The pre-trained VAE’s encoder $\mathcal{E}$ and decoder $\mathcal{D}$ are fixed during training and inference. Exclusively employed during inference, the fixed emotion predictor $\mathcal{P}$ predicts emotions on generated images for iterative emotion inference.
Our EmoEditor, in Figure 4, takes a source image $I_{s}$ and a target emotion $y$ to generate an image evoking the desired emotion. It employs the latent diffusion model (LDM) [22] with three key novelties.
@@
Integration of Global and Local Emotion Cues. In the global branch of our EmoEditor, we convert the target emotion $y$ into a binary one-hot vector $e_{oh}$, where the corresponding entry to $y$ in $e_{oh}$ is set to 1 and the rest are set to 0. This simplifies model design, enhancing comprehension and generation of complex emotional images, as one-hot encoding can capture intricate emotion combinations through a multi-label representation. We use the pre-trained emotion predictor $\mathcal{P}$ to predict the source image’s emotional state $e_{s}$, representing the probability distribution of emotions for each category. EmoEditor avoids using emotion labels for $I_{s}$ for two reasons: (1) Obtaining emotion labels for source images is often impractical in real-world scenarios. (2) Source images can be complex, conveying multiple emotions simultaneously. Utilizing a stereotypical source emotion as input potentially introduces biases and limits the model’s ability to generate diverse outputs.
@@
Next, we compute the emotion editing direction $e_{dir}$ using the target emotion code $e_{oh}$ and the source emotional state $e_{s}$: $e_{dir}=e_{oh}-e_{s}$. This allows the model to adaptively adjust the editing direction based on the initial state of different source images. To map the emotion editing direction into the latent space, we introduce an emotion encoder $\tau_{\theta}$, trained from scratch to learn neural-symbolic embeddings. It first encodes $e_{dir}$ into an emotion embedding $e$ through fully connected layers: $e=\tau_{\theta}(e_{dir})$ (See Sec S2.1 for details). Later on, we introduce novel losses to regularize the behaviors of $\tau_{\theta}$. Notably, we intentionally avoid using explicit text editing instructions as conditional inputs to the model due to the varied interpretations of target emotions, where multiple solutions can elicit the same emotions. Including explicit text instructions may hinder the model’s creativity, restricting its capacity to produce diverse sets of images conveying the same emotions.
@@
In the local branch of our EmoEditor, we introduce an image encoder $\mathcal{E}$ to extract visual features of the source image, acquiring $z$ for $I_{s}$: $z=\mathcal{E}(I_{s})$. Similar to the LDM [22], $\mathcal{E}$ is based on a pre-trained VAE [12]. The forward diffusion process gradually adds noise to $z$, generating a noisy latent $z_{t}$, where the noise level increases over $t$ time steps ($t\in T$). Specifically, we expressed $z_{t}=\sqrt{\alpha_{t}}\cdot z_{t-1}+\sqrt{1-\alpha_{t}}\cdot\epsilon$, where $\alpha_{t}$ is the diffusion coefficient controlling the rate of noise increase, and $\epsilon$ is a random noise sampled from a normal distribution.
@@
Importantly, we differentiate our alignment loss from the losses typically employed in classical language models, which focus on predicting specific word tokens. Our approach loosens the constraint of precisely predicting the exact text instructions. This flexibility empowers our model to explore a wider array of editing solutions, that can augment the source emotions of human viewers to target emotions. Consequently, our model learns to reason in a neuro-symbolic manner akin to humans and preserves its creativity and capacity for exploring novel editing solutions.
Iterative Emotion Inference (IEI). Editing emotions is a complex process often requiring iterative image edits. To address this, we introduce a recurrent emotion critic process during inference. This critic iteratively infers the evoked emotions from generated images while assessing their structural coherence and semantic consistency. Specifically, we employ the emotion predictor $\mathcal{P}$ as the critic.
In the first iteration of image generation, the source image $I_{s}$ serves as both the input and condition image, producing latent variables $z$ and $z_{t}$. These, along with the emotion vector $e$, are fed to the EmoEditor to obtain the generated image $\hat{I}$. The critic evaluates the generated image quality based on two criteria: (1) the structural similarity (SSIM) between $\hat{I}$ and $I_{s}$ falls within 0.5-0.8, and (2) the predicted emotion by $\mathcal{P}$ on $\hat{I}$ matches the target emotion, with a confidence level exceeding 0.6. If both criteria are met, EmoEditor ceases image generation; otherwise, it continues to generate new images. In subsequent iterations, we use $\hat{I}$ as the input image, generating its noisy latent variable $\hat{z}_{t}$, with $I_{s}$ remaining as the condition. We use $\mathcal{P}$ to predict the emotional state of $\hat{I}$ as the new source emotional state and calculate the new emotion editing direction to update the emotion embedding $e$. Based on these new inputs, EmoEditor generates a new edited image for the critic. This process iterates until both criteria are satisfied, or the number of critic iterations exceeds a predefined limit of 30. In the latter case, we select the generated result with the highest confidence level predicted by $\mathcal{P}$ given the target emotion among all past iterations. See Sec S2.5 for more implementation details.